Required information

Bunker makes two types of briefcase, fabric and leather. The company is currently using a traditional costing system with labor hours as the cost driver but is considering switching to an activity-based costing system. In preparation for the possible switch, Bunker has identified two activity cost pools: materials handling and setup. Pertinent data follow:

Fabric Case Leather Case
Number of labor hours 14,000 6,000
Number of material moves 504 696
Number of setups 27 63

Total estimated overhead costs are $143,400, of which $120,000 is assigned to the materials handling cost pool and $23,400 is assigned to the setup cost pool.

1. Calculate the overhead assigned to the fabric case using the traditional costing system based on direct labor hours.
2. Calculate the overhead assigned to the fabric case using ABC.
3. Was the fabric case over- or undercosted by the traditional cost system compared to ABC?
Answer #1

SOLUTION

1.

Total estimated overhead costs (A) $143,400
Total labor hours (14,000+6,000) (B) 20,000
Overhead allocation rate (C) = A/B 7.17
Total labor hours used by Fabric case (D) 14,000
Overhead assigned to the fabric case = C*D 100,380

2.

Cost Pool Total cost (A) Total Activity (B) Activity Rate (C =A/B)
Materials handling 120,000 (504+696) 1,200 100 per material move
Setup 23,400 (27+63) 90 260 per setup

Cost assigned-

Cost Pool Activity Rate (A) Activity (B) Allocated Cost (A*B)
Materials handling 100 per material move 504 50,400
Setup 260 per setup 27 7,020
Cost assigned to fabric case 57,420

3.

Overhead allocated As per Traditional Method 100,380
Overhead allocated As per ABC 57,420
Over allocated 42,960